# Load Balancing Mixture of Experts with Similarity Preserving Routers

### v1

@@ 2.2 Expert Routing
Minimizing this loss encourages the router to distribute tokens more evenly across experts. However, it may require tuning of a loss coefficient $\alpha$ to avoid overpowering the main training objective.
@@ 3.1 Load Balancing via Orthonormal Routers
This loss is dataset-agnostic and computationally cheap. We additionally initialize the router with a (near) orthogonal initialization (Saxe et al., 2014) (though it should be sufficient to simply run a few router-only training steps, see Table 2), as we find it results in quicker convergence. We call this method SimBal, as we are effectively balancing by preserving the pair-wise similarity of the tokens. We scale this loss with a coefficient of 0.1, as we find this marginally improves model perplexity. Without the coefficient, it still far outperforms previous methods.
Table 1: Parameters used for the model architecture and training. Parameter (active, total) counts include token embeddings.
@@ 3.2 Model Architecture and Training
For MoE-M and Dense-M, with $230$M active parameters and $77$M in embeddings, trained on $2\times 10^{10}$ tokens, this results in:
@@
For MoE-L and Dense-L, with $761$M active parameters and $154$M in embeddings, trained on $7.8\times 10^{10}$ tokens, this results in:
@@ 4.1 Orthogonalization and Balancing
As PyTorch currently lacks support for orthogonal parameterizations in lower-precision formats commonly used to train language models (that we use), we perform orthogonalization in float32, and then cast the resulting matrix to bfloat16, our training precision. Our loss-based method trains the matrix directly in bfloat16. We report both the maximum and mean deviation from orthonormality, as well as the final loss values, in Table 2. We find that our loss consistently produces matrices that more closely approximate orthonormality than direct orthogonal parameterizations in our scenario. In fact, our approach matches or exceeds the throughput of efficient orthogonal parameterizations, while avoiding the need for expensive reorthogonalization steps. For this synthetic experiment, we train with AdamW (with no weight decay), and a learning rate cosine decayed from $1\times 10^{-4}$ to $1\times 10^{-5}$ over 100 consecutive steps. In our MoEs, we simply add our loss as an auxiliary loss term and update once per language model training step.
In terms of expert utilization in MoEs, our method avoids collapse comparably to LBL, ensuring that no experts remain unutilized. Figure 2 illustrates the unique expert usage over time at two different scales, compared to LBL and using no losses (which results in unused experts).
@@
Figure 2: Expert utilization throughout training for MoE-M (left) and MoE-L (right), comparing LBL, our method (SimBal), and a baseline with no load balancing. We measure the number of unique experts activated on our full 77M-token validation set over time. Without any balancing, the expert routing collapses to a smaller set of experts. Both LBL and SimBal maintain full expert utilization across training. The no-loss baseline was truncated early.
@@
Figure 3: Analysis of expert redundancy in MoE-L models. (a) PES across different layers, our approach (blue) maintains significantly lower redundancy than LBL (orange). Darker = later in training. (b) Rate of change of PES during training, averaged over all layers. Redundancy occurs when many distinct experts see similar tokens, and is most likely to happen early in training, as we observe. We note that this is $>0$ at most points for LBL, suggesting it exacerbates redundancy during the majority of training.
Table 3: Model setup and performance. The best perplexity per-category is bolded.
@@
SEU $\uparrow$ 1.0 1.0 1.0 0.991
@@ 4.2 Language Modeling
We compare our method to LBL by training language models according to the setup described in Section 3.2, evaluating performance based on the perplexity of the final checkpoint. The resulting models are reported in Table 3. We additionally report Sequence-wise Expert Utilization (SEU), as reported by the mean over the fraction of experts used per sequence, to show that load balance within a sequence is not significantly degraded.
Across both MoE-M and MoE-L scales, SimBal converges approximately 36% faster than LBL. We show validation values during training in Figure 1 For MoE-L, SimBal approaches the target loss after processing roughly 50B tokens, compared to 78.6B for LBL—a 36% improvement. Similarly, in the MoE-M setting, SimBal reaches comparable loss levels at around 12.7B tokens, versus 19.9B for LBL. Overall, our method significantly outperforms LBL in both final performance and training efficiency.
@@ 4.3 Redundancy and Specialization in Experts
Motivated by the work of Dai et al. (2024), we investigate expert specialization and redundancy. As detailed in Section 3.3, we assess these properties using Pairwise Expert Similarity (PES), in contrast to their expert dropout experiments. We also do not perform expert dropping and randomized expert experiments, as we find they lack the granularity necessary to capture fine-grained redundancy patterns, and have prohibitive computational costs when comprehensively evaluating many experts. In contrast, PES offers a lightweight and scalable means of quantifying redundancy, enabling per-layer, per-checkpoint analysis across all experts in parallel.
We hypothesized that our method would lead to less redundant experts compared to LBL. This stems from the fact that LBL promotes a uniform expert distribution throughout training, which can lead to instability in the early training stages. Frequent shifts in routing distributions—caused by changing token embeddings—can induce abrupt changes in expert assignment. When the expert distribution is nearly uniform, even minor input perturbations can result in a different expert being selected. We estimate this change in expert distribution by measuring the change in redundancy, as the main source of redundancy is many experts seeing similar sets of tokens, which is exacerbated by the sensitivity of the uniform expert distribution.
As shown in Figure 3(b), the majority of expert redundancy emerges early in training, coinciding with the highest volatility in LBL (orange). During this phase, embeddings evolve rapidly, leading to unstable routing behavior and increased expert redundancy. Additionally, the change in expert redundancy is generally significantly above 0 for most of training– reinforcing our claim that LBL exacerbates redundancy.
In contrast, our method (blue) exhibits much greater stability. While expert distributions still adapt alongside evolving embeddings, they stabilize quickly. As illustrated in Figure 3(a), this leads to a consistently lower PES in the final model. Additionally, in Figure 3(b), using our loss results in a rate of change very close to zero for most of training, showing that we do not face the same redundancy-encouraging issue of LBL.
@@
4.4 Coefficient Sensitivity
Table 4: Performance across three scaling coefficients to SimBal.
@@ 4.4 Coefficient Sensitivity
We examine the coefficient sensitivity of SimBal to determine if tuning is necessary. We test coefficients 0.01, 0.1, and 1.0, and show our results in Table 4. We observe that there is slight benefit, but it is not very sensitive past a certain point. We recommend using a coefficient 0.1 for most setups. We additionally find that the final min PES values are lower across the board compared to using LBL.
(a) MoE-L
@@ 5 Related Work
Traditional approaches employ an auxiliary load balancing loss (Fedus et al., 2022) to encourage a uniform distribution over experts, which can interfere with the main training objective and potentially degrade performance. To address this, auxiliary-loss-free (Wang et al., 2024) strategies have been introduced, notably used in DeepSeek-V3 (DeepSeek-AI et al., 2025), but in conjunction with auxiliary balancing loss. This method dynamically adjusts per-expert bias terms added to the routing scores, guiding the top-K expert selection without introducing additional gradients. While effective at improving global load balance, it struggles in balancing the MoE sequence-wise, which can cause degradations in throughput. We explore improving loss-free load balancing using our auxiliary loss in Appendix A.1. Additionally, LF is sensitive to batch size, with (Wang et al., 2024) finding a 0.5 decrease in perplexity when using a batch size of 512 rather than 4 (per-device, no sync), a property that is much less drastic in LBL and irrelevant with SimBal (as it is invariant to the data). (Wang et al., 2024) also claims LBL achieves improved specialization with a distributed sync to maximize batch size, we eliminate the need for this entirely.
@@ 6 Limitations
While we train our models with relatively large data multipliers, prior work such as Muennighoff et al. (2025) suggests that substantially more data may be necessary to achieve strong performance on downstream benchmarks. Nevertheless, our training setup provides sufficient scale to meaningfully compare the relative effectiveness of different balancing methods.
Finally, although our architectural choices align with recent MoE literature, our study is limited to a single set of design decisions. We leave the exploration of alternative configurations to future work. For instance, we do not investigate how token dropping might affect the performance of our balancing mechanism (focusing on higher-quality dropless models (Gale et al., 2022)), which could be a valuable direction for further analysis.
@@ 7 Conclusion
In this work, we introduced a novel load balancing mechanism for Mixture-of-Experts (MoE) models that consistently outperforms popular approaches across multiple scales. We also proposed efficient, scalable metrics for quantifying expert redundancy, and demonstrated that models with lower redundancy—as measured by our proposed metric—exhibit improved parameter efficiency.
Appendix A Supplementary Material
@@ A.1 Loss-Free Load Balancing Combination
Table 5: Model setup and performance. Sequence-wise Expert Utilization (SEU) is computed as the mean over the fraction of activated experts within a sequence. SimBal can improve sequence-wise balance without significant performance degradation, sometimes improving performance. All models use all experts throughout the full validation set, LF is the least balanced per-batch. While LBL asserts near-perfect balance, it also causes substantial degradation.
@@
Other work [DeepSeek-AI et al., 2025] use LBL in conjunction with LF for batch-wise load balancing, as they find that it can result in substantial imbalance in expert use sequence-wise. We do not include these results in the main paper due to this imbalance and requirement for additional losses. In this section, we explore whether a combination with SimBalworks similarly to LBL to improve sequence-wise balancing.
While the original authors of LF use sigmoid gating (over our softmax gating), we find that softmax gating is substantially more common in state-of-the-art work. Thus, to maximize relevance (regardless of performance), we compare using softmax in the main paper. Regardless, we still include sigmoid LF comparisons in this section. The training setup for MoE-M remains identical to Section 3.2 otherwise. We find that sigmoid gating resulted in substantially worse sequence-wise balance.
We evaluated the balancing capabilities of this method using the MoE-M configuration, comparing its performance against both LBL and SimBal. We also hypothesize that combining LF balancing with SimBal would result in improved performance for our method, as the bias term learned for routing was very similar in LBL and SimBal, but different from the LF bias. We summarize our results in Table 5. We find that sigmoid gating improves perplexity, but leads to significant degradation in sequence-wise balance, especially compared to using only SimBal or LBL (as seen in Table 3). This is not ideal, as with larger models, when using model parallelism, extra consideration may be needed to ensure full utilization of all devices. Using LBL mitigates some of this, but leads to a substantial degradation in performance, thus indicating that SimBal may strike a good balance between batch-wise balance and model perplexity.

### v2

@@ 2.2 Expert Routing
Minimizing this loss encourages the router to distribute tokens more evenly across experts. However, it may require tuning of a loss coefficient $\alpha$ to avoid overpowering the main training objective. We include PyTorch implementation details in Appendix A.3.
@@ 3.1 Load Balancing via Orthonormal Routers
This loss is dataset-agnostic and computationally cheap. This is important, as Qiu et al. (2025) finds that existing losses, which are dependent on the data, require large batch sizes to be effective. We additionally initialize the router with a (near) orthogonal initialization (Saxe et al., 2014) (though it should be sufficient to simply run a few router-only training steps, see Table 2), as we find it results in quicker convergence. We call this method SimBal, as we are effectively balancing by preserving the pair-wise similarity of the tokens. The experiments in our paper scale this coefficient by 0.1, but we find that this is not important, as shown in Section 4.2. We include PyTorch implementation details in Appendix A.3.
Table 1: Parameters used for the model architecture and training. Parameter (active, total) counts include token embeddings. All MoE models have 32 experts, with the top 4 activated.
@@ 3.2 Model Architecture and Training
For MoE-M and Dense-M, with 230230M active parameters and 7777M in embeddings, trained on $2\times 10^{10}$ tokens, this results in:
@@
For MoE-L and Dense-L, with 761761M active parameters and 154154M in embeddings, trained on $7.8\times 10^{10}$ tokens, this results in:
@@ 4.1 Orthogonalization and Balancing
As PyTorch currently lacks support for orthogonal parameterizations in lower-precision formats commonly used to train language models (that we use), we perform orthogonalization in float32, and then cast the resulting matrix to bfloat16, our training precision. Our loss-based method trains the matrix directly in bfloat16. We report both the maximum and mean deviation from orthonormality, as well as the final loss values, in Table 2. We find that our loss consistently produces matrices that more closely approximate orthonormality than direct orthogonal parameterizations in our scenario. In fact, our approach matches or exceeds the throughput of efficient orthogonal parameterizations, while avoiding the need for expensive reorthogonalization steps. For this synthetic experiment, we train with AdamW (with no weight decay), and a learning rate cosine decayed from $1\times 10^{-4}$ to $1\times 10^{-5}$ over 100 consecutive steps. In our MoEs, we simply add our loss as an auxiliary loss term and update once per language model training step. We examine the coefficient sensitivity of SimBal to determine if tuning is necessary.
In terms of expert utilization in MoEs, our method avoids collapse comparably to LBL, ensuring that no experts remain unutilized. Figure 2 illustrates the unique expert usage over time at two different scales, compared to LBL and using no losses (which results in unused experts). To verify that sequence-wise balance is not substantially degraded, we compare SimBal against LBL by measuring the entropy of the routing distributions and Sequence-wise Expert Utilization (SEU), as reported by the mean over the fraction of experts used per sequence, to show that load balance within a sequence is not significantly degraded. We report our results in Table 3.
To analyze whether SimBal is able to effectively orthogonalize routing matrices, we analyze the mean layer-wise L2 distance of the final router gram matrix from the identity matrix in Table 3. More in-depth data with layer-wise values across MoE-L and MoE-M can be found in Appendix A.2.
Table 3: Load balancing and orthogonalization of LBL and SimBal on MoE-L.
Metric SEU Entropy $(R^{T}R-I)^{2}$
LBL 1.000 1.268 0.0311
SimBal 0.991 1.168 2.121 $\times$ 10-8
@@
Figure 2: Expert utilization throughout training for MoE-M (left) and MoE-L (right), comparing LBL, our method (SimBal), and a baseline with no load balancing. We measure the number of unique experts activated on our full 77M-token validation set over time. Without any balancing, the expert routing collapses to a smaller set of experts. Both LBL and SimBal maintain full expert avoid expert collapse. The no-loss baseline was truncated early.
@@
Figure 3: Analysis of expert redundancy in MoE-L models. (a) PES across different layers, our approach (blue) maintains significantly lower redundancy than LBL (orange). Darker = later in training. (b) Rate of change of PES during training, averaged over all layers. Redundancy occurs when many distinct experts see similar tokens, and is most likely to happen early in training, as we observe. We note that this is >0>0 at most points for LBL, suggesting it exacerbates redundancy during the majority of training.
Table 4: Model setup and performance.
@@
Table 5: Comparison of LBL-L and SimBal-L performance across benchmarks.
Benchmark LBL-L $\pm$ stderr SimBal-L $\pm$ stderr
ARC Challenge (Clark et al., 2018) $22.44\%\pm 1.22\%$ $23.21\%\pm 1.23\%$
ARC Easy (Clark et al., 2018) $40.49\%\pm 1.01\%$ $41.16\%\pm 1.01\%$
HellaSwag (Zellers et al., 2019) $35.45\%\pm 0.48\%$ $35.74\%\pm 0.48\%$
PIAQ (Bisk et al., 2019) $6.49\%\pm 1.10\% 66$$6.81\%\pm 1.10\%$
WinoGrande (Sakaguchi et al., 2019) $49.72\%\pm 1.41\%$ $52.49\%\pm 1.40\%$
GLUE (Wang et al., 2018) $45.10\%\pm 1.98\%$ $51.73\%\pm 1.97\%$
mean $43.28\%$ $45.19\%$
@@ 4.2 Language Modeling
We compare our method to LBL by training language models according to the setup described in Section 3.2, evaluating performance based on the perplexity of the final checkpoint. The resulting models are reported in Table 4. We additionally report the SEU of the models.
Across both MoE-M and MoE-L scales, SimBal converges approximately 36% faster than LBL. We show validation values during training in Figure 1 For MoE-L, SimBal approaches the target loss after processing roughly 50B tokens, compared to 78.6B for LBL—a 36% improvement. Similarly, in the MoE-M setting, SimBal reaches comparable loss levels at around 12.7B tokens, versus 19.9B for LBL. We additionally evaluate MoE-L on standard downstream benchmarks to test whether the perplexity gains of SimBal translate to broader tasks, comparing against LBL (Table 5). Overall, our method outperforms LBL in both downstream performance and training efficiency.
We train 4 additional models (for a total of 5 models) for both SimBal and LBL on MoE-M (due to computational limitations) to parse the statistical significance of our results. We find that models trained with LBL have a mean perplexity of 14.051 with a standard deviation of 0.026. In comparison, SimBal achieves a mean perplexity of 13.691 with standard deviation 0.039. The mean SimBal performance is over 13 standard deviations lower than the perplexity of LBL, showing that our results are very statistically significant.
Finally, we examine sensitivity to the auxiliary loss coefficient (0.01, 0.1, 1.0), with results in Table 7. Based on our 5-seed runs on MoE-M, the effect is negligible, and we do not recommend tuning this hyperparameter.
@@ 4.3 Redundancy and Specialization in Experts
(a) MoE-L
Figure 4: Number of dropped top experts vs. validation loss, as proposed by Dai et al. (2024). SimBal exhibits lower redundancy, as shown by worse performance as more experts are dropped.
(a) MoE-L
Motivated by Dai et al. (2024), we study expert specialization and redundancy. As described in Section 3.3, we measure these properties with Pairwise Expert Similarity (PES), in contrast to their expert dropout approach. In Figure 5, we validate the correlation between PES and their method, reproducing their redundancy analysis. By their metric, SimBal shows lower redundancy, as validation perplexity rises more sharply when top experts are dropped. However, such dropout-based metrics lack granularity and are prohibitively expensive for large-scale evaluation. PES instead provides a lightweight, scalable measure of redundancy, enabling per-layer, per-checkpoint analysis across all experts in parallel.
We hypothesize that SimBal produces less redundant experts than LBL. LBL enforces uniform distributions, leading to instability in early training as changing embeddings cause frequent routing shifts. Under near-uniform assignment, small input perturbations can reassign tokens, creating redundancy as many experts see similar tokens. We capture this effect by measuring changes in redundancy.
As shown in Figure 3(b), most redundancy in LBL (orange) arises early, coinciding with embedding volatility and unstable routing. Redundancy remains above zero through much of training, reinforcing that LBL amplifies it. In contrast, SimBal (blue) stabilizes quickly: while expert distributions adapt, they converge to consistently lower PES (Figure 3(a)). Moreover, the rate of change remains near zero for most of training, showing that our method avoids the issues of LBL.
@@
4.4 Inference-Time Expert Pruning
Table 6: Dynamic-K expert stage 3 selection (Szatkowski et al., 2024) synergy with SimBal vs. LBL (perplexity and runtime on a full validation run, MoE-L). SimBal is able to preserve more performance when lower weighted experts are dropped for faster runtime.
Dropped $P(E)$ < SimBal (PPL) SimBal (s) LBL (PPL) LBL (s)
0 8.317 620.927 8.536 619.657
0.1 8.364 571.147 8.542 575.121
0.15 8.598 503.065 8.621 543.027
0.2 9.380 472.915 9.057 495.200
We further evaluate SimBal under inference-time expert pruning, following Szatkowski et al. (2024), where experts with assignment probabilities below a threshold are dropped at runtime. Results are presented in Table 6. SimBal produces less uniform assignments, allowing pruning to drastically improve efficiency with minimal perplexity cost. In contrast, LBL shows weaker synergy with pruning: while its performance drop is smaller (likely due to redundancy, similarly to Figure 4(a)), improvements in throughput are limited. Notably, when experts below a weight of $0.15$ are dropped (where both perplexities are most similar), SimBal achieves a 7.4% speedup (543s vs. 503s).
Table 7: Performance across three scaling coefficients to SimBal. We find that the differences are not significant enough to warrant hyperparameter tuning.
@@ 5 Related Work
Traditional approaches employ an auxiliary load balancing loss (Fedus et al., 2022) to encourage a uniform distribution over experts, which can interfere with the main training objective and potentially degrade performance. To address this, auxiliary-loss-free (LF) strategies have been introduced (Wang et al., 2024), notably used in DeepSeek-V3 (DeepSeek-AI et al., 2025), but always in conjunction with an auxiliary balancing loss. LF dynamically adjusts per-expert bias terms added to the routing scores, guiding top-$K$ expert selection without introducing additional gradients. While this improves global balance, it struggles to balance MoE usage sequence-wise, often degrading throughput.
Due to difficulties in achieving effective load balance in our early experiments, we did not pursue full-scale MoE-L training with LF in the main paper, and instead provide an in-depth analysis in Appendix A.1. Moreover, LF is highly sensitive to batch size: Qiu et al. (2025) report a substantial perplexity drop when training with batch size $512$ vs. $4$ (per-device, no sync). This effect is far milder for LBL, and entirely absent for SimBal, which is invariant to the data. Finally, while Qiu et al. (2025) argue that LBL requires distributed synchronization to maximize batch size and improve specialization, SimBal eliminates this need altogether.
@@ 6 Limitations
While we train our models with relatively large data multipliers, prior work such as Muennighoff et al. (2025) suggests that substantially more data (trillions of tokens) may be necessary to achieve strong performance on downstream benchmarks. Nevertheless, our training setup provides sufficient scale to meaningfully compare the relative effectiveness of different balancing methods, which we supplement with statistical significance comparisons.
Finally, although our architectural choices align with recent MoE literature, our study is limited to a single set of design decisions. We leave the exploration of alternative configurations to future work. For instance, we do not investigate how token dropping might affect the performance of our balancing mechanism (instead focusing on higher-quality dropless models (Gale et al., 2022)), which could be a valuable direction for further analysis.
@@ 7 Conclusion
In this work, we introduced a novel load balancing mechanism for Mixture-of-Experts (MoE) models that consistently outperforms popular approaches across two scales. We also proposed efficient, scalable metrics for quantifying expert redundancy, and demonstrated that models with lower redundancy—as measured by our proposed metric and existing methods—exhibit improved parameter efficiency.
Appendix A Appendix
@@ A.1 Loss-Free Load Balancing Combination
Table 8: Model setup and performance. Sequence-wise Expert Utilization (SEU) is computed as the mean over the fraction of activated experts within a sequence. SimBal can improve sequence-wise balance without significant performance degradation, sometimes improving performance. All models use all experts throughout the full validation set, LF is the least balanced per-batch. While LBL asserts near-perfect balance, it also causes substantial perplexity degradation.
@@
Other work [DeepSeek-AI et al., 2025] use LBL in conjunction with LF for batch-wise load balancing, as they find that it can result in substantial imbalance in expert use sequence-wise. We do not include these results in earlier charts due to this extreme imbalance. Instead, in this section, we explore whether a combination with SimBal works similarly to LBL to improve sequence-wise balancing.
While the original authors of LF use sigmoid gating (over our softmax gating), we find that softmax gating is substantially more common in state-of-the-art work. Thus, to maximize relevance (regardless of performance), we additionally compare with softmax gating. The training setup for MoE-M remains identical to Section 3.2 otherwise.
We evaluated the balancing capabilities of this method using the MoE-M configuration, comparing its performance against both LBL and SimBal. We summarize our results in Table 8. We find that sigmoid gating leads to significant degradation in sequence-wise balance, especially compared to using only SimBal or LBL (as seen in Table 4). In exchange, there was a minor and possibly statistically insignificant (using the deviation values from Section 4.2. This is not ideal, as with larger models, when using model parallelism, extra consideration may be needed to ensure full utilization of all devices. Using LBL mitigates some of this, but leads to a substantial degradation in performance.
A.2 Layer-Wise Orthogonalization
We provide tables for layer-wise orthogonalization performance for SimBal, and compare the results to LBL on MoE-M (Table 9) and MoE-L (Table 10). LBL alone does not orthogonalize the router whatsoever, while SimBal is able to achieve mean squared error similar to commonly used $\epsilon$ for numerical stability.
Router SimBal LBL
Layer 0 Router 1.94017e-10 0.00146701
Layer 1 Router 1.70156e-10 0.01486
Layer 2 Router 1.91267e-10 0.0155954
Layer 3 Router 1.89254e-10 0.0102319
Layer 4 Router 1.50925e-08 0.0100937
Layer 5 Router 2.99727e-08 0.0143029
Layer 6 Router 1.82301e-10 0.020765
Layer 7 Router 1.73648e-10 0.0258847
Table 9: Router orthogonality of MoE-M, as measured by $(R^{T}R-I)^{2}$
Router SimBal LBL
Layer 0 Router 1.49951e-08 0.0125956
Layer 1 Router 1.00854e-10 0.027788
Layer 2 Router 1.03228e-10 0.0183506
Layer 3 Router 4.47955e-08 0.0128958
Layer 4 Router 1.5001e-08 0.00668315
Layer 5 Router 9.38376e-11 0.00399825
Layer 6 Router 1.16159e-10 0.00375414
Layer 7 Router 2.99078e-08 0.00736187
Layer 8 Router 4.47949e-08 0.0200508
Layer 9 Router 2.99088e-08 0.0377724
Layer 10 Router 5.97087e-08 0.083971
Layer 11 Router 1.49907e-08 0.138501
Table 10: Router orthogonality of MoE-L, as measured by $(R^{T}R-I)^{2}$
A.3 Implementation Details
Here we provide some implementation details related to the auxiliary losses used in the paper in Figure 6. For our LBL baseline, we use an open-source repository implementation based on Zoph et al. [2022], available at lucidrains/st-moe-pytorch on GitHub. For both, we multiply the output of the function by the scaling coefficient if/where applicable during training. These losses can then be added to the final model loss (by adding them), or included using the AddAuxiliaryLoss autograd trick used in DeepSeek’s modeling_deepseek.py on HuggingFace.
1import torch
2from einops import reduce
3
4# LBL
5def balance_loss(gates: torch.Tensor) -> torch.Tensor:
6 batch_size, num_tokens, num_experts = gates.shape
7
8 # bal_loss = E * sum(f_i * P_i), expert i
9 expert_mask = gates > 0.0
10 f_i = reduce(expert_mask.float(), "b t e -> b e", "mean")
11 P_i = reduce(gates, "b t e -> b e", "mean")
12 loss_per_batch = num_experts * torch.sum(f_i * P_i, dim=-1)
13 return loss_per_batch.mean()
14
15# SimBal
16def simbal_loss(router_linear, p=1):
17 w = router_linear.weight
18 w_ortho = torch.matmul(w, w.T)
19 eye = torch.eye(w.shape[0], device=w.device)
20 loss = torch.norm(w_ortho - eye, p=p)
21 return loss
Figure 6: Python implementations of the LBL and SimBal loss functions.